3. Given a frustum of a right circular cone with a slant height of 9ft ., and the radii of bases are 5ft and 7ft. Find:







a. The lateral area







b. Altitude of the frustum







c. The altitude of the entire cone if the cone removed was replaced








1
Expert's answer
2021-12-07T09:36:03-0500

Given,

Slant height of right circular cone "(l_1)=9ft"

Base radius "(r_1)= 5ft"

and "(r_2)=7ft"

a) Lateral area of the frustum "=\pi (r_2-r_1)l_1"

"=3.14 (7-5)\times 9 \\
=3.14\times 2\times 9 \\
=59.52ft^2"

b) Altitude of the frustum

"h'=(\sqrt{9^2-4^2})=\sqrt{77}=8.87ft"

"\cos(\theta)=\frac{r_2-r_1}{l_1} \\
=\frac{7-2}{9} \\
=\frac{2}{9}\\
\Rightarrow \theta = \cos^{-1}(\frac{2}{9}) \\
\Rightarrow \theta = 77.16^\circ"

c) Altitude of the cone complete cone

So, "\tan(\theta)=\frac{h}{7}"

"h=\tan(77.16)\times 7 = 30.7ft"